# On the Entanglement Entropy Distribution of a Hybrid Quantum Circuit

### 31 Mar 2026 v1

@@ IV.3 Numerical simulation
Figure 11: Entanglement entropy distributions in the volume law regime (right) and area law regime (left). Distributions predicted from the DPRE model (dashed line) and the FP solution (solid line) capture the shape of the distribution. The entanglement entropy scale on the horizontal axis has been shifted so that the two distributions are visible together.
In Fig. 11, we show a comparison between the histogram of our data and theoretical predictions as discussed above. The DPRE model accurately captures the distribution, whereas the solution to the FP equation captures it with moderate accuracy.
To obtain the theoretical distributions in the volume law regime, the non-universal constants $s_{0}$ and $s_{1}$ are required. From the Eq. (11), they can be found simultaneously from the $\langle S\rangle$ and $\mathrm{Var}[S]$ for a given $p$:
@@
In Fig. 10, we compare the IoD from our data with that predicted by the DPRE model for $p=0.08$. From the figure, we see that the IoD for different system sizes follows the computed trend (dashed line).
Apart from $\mu$ and $\nu$, which are the mean and variance of the change in entanglement entropy induced by a single unitary layer, the parameter $c$ is required to obtain the resulting distribution in the area-law regime. From our simple classical model, it was expected to scale $q^{2}$. However, we found that such scaling of $c$ does not yield an accurate estimate so we fit $c$ with a generic quadratic function, obtaining $c=0.3285q^{2}-0.4536q+0.1239$.
Figure 12: Comparison of IoD (top) and skewness (bottom) curves between the solution to the FP equations and the numerical data in the area law regime.
In Fig. 12, we compute the IoD and skewness from the predicted distributions in the area-law regime. In terms of IoD, it can be seen that there is a large gap between the IoD predicted from the model and data, which closes in the large $p$ limit. The skewness curve of the predicted distribution follows a similar trend to that of the fitted curve and the data.
Figure 13: Kullback-Leibler divergence between data and predicted model.
Finally, to assess how accurately the predicted distributions describe the numerical data across a wide range of measurement rates, we compute the Kullback–Leibler (KL) divergence [26] between the model prediction, $Q(S)$ and the simulated distributions, $P(S)$, (Fig. 13) where the KL divergence is defined by:
$D_{\mathrm{KL}}(P||Q)=\sum_{s\in S}P(s)\log\frac{P(s)}{Q(s)}$ (30)
where the sum over $s$ runs over all realizations of entanglement entropy trajectories with non-zero values of $P(S)$. If the $P$ and $Q$ are exactly the same, $D_{\mathrm{KL}}=0$.
The figure indicates that both models fail at criticality. Nevertheless, we find that the TW distribution derived from the DPRE mapping accurately captures the data in the volume-law regime, whereas the FP solution provides an excellent description in the area-law regime.

### 13 Apr 2026 v2

@@ IV.3 Numerical simulation
Figure 11: Entanglement entropy distributions in the volume law regime (right) and area-law regime (left). Distributions predicted by the DPRE model (dashed line) and the Fokker-Planck equation (solid line) capture the shape of the distribution. The entanglement entropy scale on the horizontal axis has been shifted so that the two distributions are visible together.
We obtain the theoretical distributions in the volume-law regime by fixing the non-universal constants $s_{0}$ and $s_{1}$. From the Eq. (11), they can be found simultaneously from the $\langle S\rangle$ and $\mathrm{Var}[S]$ for a given $p$:
@@
To obtain the solution of the Fokker-Planck equation, the parameters $\mu$, $\nu$, and the function $c(q)$ must be specified. The mean and variance of the entanglement entropy change induced by a single unitary layer, $\mu$ and $\nu$, are extracted numerically by tracking $\Delta S$ under one unitary layer, which we find to be $\mu=0.58$ and $\nu=0.80$. The covariance correction $c(q)$ is then determined by fitting the numerical data to a quadratic function of $q$, which gives $c(q)=0.3285q^{2}-0.4536q+0.1239$.
In Fig. 11, we show a comparison between the histogram of our data and theoretical predictions as discussed above. The DPRE model accurately captures the distribution, whereas the solution of the Fokker-Planck equation captures it with moderate accuracy.
Figure 12: Comparison of IoD (top) and skewness (bottom) curves between the solution of the Fokker-Planck (FP) equations and the numerical data in the area-law regime.
In Fig. 12, we compute the IoD and skewness from the predicted distributions in the area-law regime. In terms of IoD, there is a large gap between the IoD predicted by the model and the simulated data, which closes in the large-$p$ limit. The skewness curve of the predicted distribution follows a similar trend to that of the fitted curve and the data.
Figure 13: Kullback-Leibler divergence between the simulation data and the effective models derived from DPRE and the Fokker-Planck equation.
Finally, to assess how accurately the predicted distributions describe the numerical data across a wide range of measurement rates, we compute the Kullback–Leibler (KL) divergence [27] between the model prediction, $Q(S)$, and the simulated distributions, $P(S)$, where the KL divergence is defined as
$D_{\mathrm{KL}}(P||Q)=\sum_{s\in S}P(s)\log\frac{P(s)}{Q(s)}.$ (29)
Here, the sum over $s$ runs over all realizations of entanglement entropy trajectories with non-zero values of $P(S)$. The KL-divergence vanishes if and only if the two distributions are the same. Figure 13 shows that the TW distribution derived from the DPRE mapping and the solution of the Fokker-Planck equation provide an accurate description of the entanglement entropy distributions in the volume-law and the area-law phases, respectively, whereas both models become inaccurate as the measurement rate gets closer to criticality.
@@ Acknowledgements.
Appendix A Perturbative expansion of $P(S)$
We provide a Maclaurin expansion of the stationary distribution $P(S)$ of the Fokker-Planck equation in Eq. (26) as follows [45, 6]:
$P(S)\propto\frac{1}{b}P_{0}P_{1}P_{2},$ (30)
where $P_{0},P_{1}$ and $P_{2}$ collect the contributions of order $c(q)]^{0},[c(q)]^{1}$, and $c(q)]^{2}$. Each term can be expressed as
$P_{0}$ $=x^{\gamma}e^{-\frac{2S}{1-q}}$ (31)
$P_{1}$ $=x^{d_{1}c(q)}\exp[-c(q)\{-d_{0}x^{-1}+d_{2}x+\frac{d_{3}}{2}x^{2}\}]$ (32)
$P_{2}$ $=x^{e_{1}[c(q)]^{2}}$ (33)
$\quad\times\exp[-[c(q)]^{2}\{-e_{0}x^{-1}+e_{2}x+\frac{e_{3}}{2}x^{2}+\frac{e_{4}}{3}x^{3}\}],$ (34)
where $x=(\nu+\alpha S)$ and $\gamma=\frac{2}{\alpha}(\frac{\nu}{1-q}+\mu)$. The coefficients $d_{i}$ and $e_{i}$ are determined by expanding in $c(q)$ and then integrating over $S$:
$d_{0}$ $=2\left[\mu\left(\frac{\nu}{\alpha}\right)+(\mu+q)\left(\frac{\nu}{\alpha}\right)^{2}-q\left(\frac{\nu}{\alpha}\right)^{3}\right]$ (35)
$d_{1}$ $=2\left[\frac{3q\nu}{\alpha^{3}}-\frac{2\nu(\mu+q)}{\alpha^{2}}-\frac{\mu}{\alpha}\right]$ (36)
$d_{2}$ $=2\left[\frac{\mu+q}{\alpha^{2}}-3\frac{q\nu}{\alpha^{3}}\right]$ (37)
$d_{3}$ $=2\frac{q}{\alpha^{3}},$ (38)
and
$e_{0}$ $=-2\left[\frac{\mu}{\alpha}\nu+\frac{2\mu+q}{\alpha^{2}}\nu^{2}+\frac{\mu+2q}{\alpha^{3}}\nu^{3}+\frac{q}{\alpha^{4}}\nu^{4}\right]$ (39)
$e_{1}$ $=2\left[\frac{\mu}{\alpha}+2\left(\frac{2\mu+q}{\alpha^{2}}\right)\nu+3\left(\frac{\mu+2q}{\alpha^{3}}\right)\nu^{2}+4\left(\frac{q}{\alpha^{4}}\right)\nu^{3}\right]$ (40)
$e_{2}$ $=-2\left[\frac{2\mu+q}{\alpha^{2}}+3\left(\frac{\mu+2q}{\alpha^{3}}\right)\nu+6\left(\frac{q}{\alpha^{4}}\right)\nu^{2}\right]$ (41)
$e_{3}$ $=2\left[\frac{\mu+2q}{\alpha^{3}}+4\left(\frac{q}{\alpha^{4}}\right)\nu\right]$ (42)
$e_{4}$ $=\frac{-2q}{\alpha^{4}}.$ (43)